Hi, could you please help me with an algebra equation?
It is:
Find the value of x: -20-x=-12

If you can find the time to answer it, please do.
Thankyou!!!!!
Shannon

Answer by nerdybill(7384)

Find the value of x:
-20-x=-12
.
The idea is to isolate the unknown (in this case, 'x') to one side of the equation (doesn't matter which side).
.
-20-x=-12
Add x to both sides:
-20-x+x = -12+x
-20 = -12+x
Add 12 to both sides:
-20+12 = -12+x+12
-8 = x